A fund earning 4% effective has a balance of $1,000 at the beginning of the year. If $200 is added to the fund at the end of three months and if $300 is withdrawn from the fund at the end of nine months, find the ending balance under the assumption that simple interest approximation is used.

please solve it without table and excal
Answer #1

Answer: $943

After end of 1 year the ending amount = ending principal + simpel itnerest earned during the year.

Ending principal = 1000 + 200 - 300 = 900

Simple interest earned = 1000*4% + 200*3% - 300*1% = 40 + 6 - 3 = 43

Ending balance = 900 + 43 = $943